What is the resistance R of a 51.7 m long aluminum wire that has a diameter of 8.93mm? The resistivity of aluminum is 2.83 x 10-8


1
Expert's answer
2022-03-27T10:56:08-0400

The resistance of the wire is

"R=\rho\frac{l} {S} =\rho\frac{4l} {\pi d^2},"

where "\rho" - resistivity of aluminium,

l - length of wire,

S - cross-sectional area of wire,

d - diameter of wire.


So,


"R=2.83\cdot 10^{-8}\frac{4\cdot 51.7} {3.14\cdot (8.93\cdot 10^{-3})^2}=0.02\space Ohm."